40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 18 20:21:33 2019, shruti, Update, ALS, AM measurement attempt at X end 
    Reply  Thu Nov 7 17:46:10 2019, shruti, Update, ALS, AM measurement at X end AMTF20191107.png
Message ID: 15020     Entry time: Thu Nov 7 17:46:10 2019     In reply to: 14979
Author: shruti 
Type: Update 
Category: ALS 
Subject: AM measurement at X end 

Some details:

- There was a SR560+SR785 (not connected for measurement) placed near the X end which I moved; it is now behind the electronics rack by the X arm beam tube (~15m away).

- Also, for the AM measurement I moved the AG5395A from behind the PSL setup to the X end, where it now is.

- By toggling the XGREEN shutter, I noticed that the cavity was not resonant before I disconnected anything from the setup since the spot shape kept changing, but I proceeded anyway. 

- Because Rana said that it was important for me to mention: the ~5 USD blue-yellow crocs (that I now use) work fine for me.

The AM Measurement:

1. The cables were calibrated with the DC block in the A port (for a A/R measurement)

2. The cable to the PZT was disconnected from the pomona box and connected to the RF out of the NA, the PD output labelled 'GREEN_REFL' was also disconnected and connected to the B port via a DC block. 

3. The ITMX was 'misaligned'. (This allowed the reflected green PD output as seen on the oscilloscope to stabilize.)

4. The PZT is modulated in frequency and the residual amplitude modulation (as observed in the measured reflected green light) is plotted, ref. Attachment 1. The parameters for the plotted data in the attachment were:

# AG4395A Measurement - Timestamp: Nov 07 2019 - 17:04:07
#---------- Measurement Parameters ------------
# Start Frequency (Hz): 10000.0, 10000.0
# Stop Frequency (Hz): 10000000.0, 10000000.0
# Frequency Points: 801, 801
# Measurement Format: LOGM, PHAS
# Measuremed Input: AR, AR
#---------- Analyzer Settings ----------
# Number of Averages: 8
# Auto Bandwidth: On, On
# IF Bandwidth: 300.0, 300.0
# Input Attenuators (R,A,B): 0dB 10dB 20dB 
# Excitation amplitude = -10.0dBm

 

 

------------------------------------

Update (19:13 7thNov19):  When the ITMX was intentionally misaligned, Rana and I checked to see if the Oplevs were turned off and they were. But while I was casually checking the Oplevs again, they were on! 

Not sure what to do about this or what caused it. 

Quote:

[Shruti, Rana]

- At the X end, we set up the network analyzer to begin measurement of the AM transfer function by actuation of the laser PZT.

- The lid of the PDH optics setup was removed to make some checks and then replaced.

- From the PDH servo electronics setup the 'GREEN_REFL' and 'TO AUX-X LASER PZT' cables were removed for the measurement and then re-attached after.

- The signal today was too low to make a real measurement of the AM transfer function, but the GPIB scripts and interfacing was tested. 

 

Attachment 1: AMTF20191107.png  110 kB  | Hide | Hide all
AMTF20191107.png
ELOG V3.1.3-